# Local Normalization Distortion and the Thermodynamic Formalism of Decoding Strategies for Large Language Models

### v1

@@ Abstract
Advances in hardware and language model architecture have spurred a revolution in natural language generation. However, autoregressive models compute probability distributions over next-token choices, and sampling from these distributions, known as decoding, has received significantly less attention than other design choices. Existing decoding strategies are largely based on heuristics, resulting in methods that are hard to apply or improve in a principled manner. We develop the theory of decoding strategies for language models by expressing popular decoding algorithms as equilibrium states in the language of ergodic theory and stating the functions they optimize. Using this, we analyze the effect of the local normalization step of top-k, nucleus, and temperature sampling, used to make probabilities sum to one. We argue that local normalization distortion is a fundamental defect of decoding strategies and quantify the size of this distortion and its effect on mathematical proxies for the quality and diversity of generated text. Contrary to the prevailing explanation, we argue that the major cause of the under-performance of top-k sampling relative to nucleus sampling is local normalization distortion. This yields conclusions for the future design of decoding algorithms and the detection of machine-generated text.
@@ 1 Introduction
Autoregressive large-language models are poised to transform industries such as healthcare, finance and education Zhao et al. (2024). Rapid advancements in this field have been fueled by scaling Hoffmann et al. (2024), transformer network architectures Vaswani et al. (2017), and alignment processes Christiano et al. (2017); Rafailov et al. (2024). However, autoregressive language models produce conditional distributions over predicted next tokens, and these must be iteratively sampled during inference. A suite of decoding methods exist for this sampling, such as top k, nucleus, or temperature sampling, but these are based on heuristics and our understanding of these methods is in its infancy, with most being developed through trial and error Wiher et al. (2022).
This is surprising since the choice of decoding strategy has a profound impact on the quality of the generated text and, in some cases, may be more important than the model architecture Wiher et al. (2022). For example, greedy sampling, or the related concept of beam search, tends to produce accurate but repetitive or dull text Holtzman et al. (2020). In contrast, pure sampling produces much more varied and interesting text but can produce text which is ‘incoherent and almost unrelated to the context’ Holtzman et al. (2020).
Our primary goal is to fill this gap and initiate the theoretical study of decoding strategies. As an application of this theory, we investigate a well-known defect of popular decoding strategies, which results from truncating the distribution of possible next tokens at each step during inference. Truncation necessitates repeated renormalization to obtain valid probability distributions, and results in a phenomenon we term local normalization distortion. The fact that local normalization distorts the resulting probability distribution is well-known, but not well-understood. Local normalization distortion measures deviation from the learned conditional distributions of a language model, and we theoretically and empirically validate claims on its relationship to the quality of generated text. In summary, our contributions are as follows.
@@
    Develop a theoretical framework for analyzing decoding strategies. This involves describing the probability distributions $q$ produced by decoding algorithms as equilibrium states, drawing heavily on the language of ergodic theory and thermodynamic formalism Bowen (2008). We precisely state the function maximized by each decoding strategy; see Section 5.
@@
    Evaluate language models and decoding strategies in terms of a quality-diversity trade-off, as in Caccia et al. (2019). We show, both theoretically and empirically, that local normalization distortion negatively affects the performance of the decoding strategy. Indeed, we show that the widely noted quality-diversity under-performance of top-k sampling relative to nucleus (top-p) sampling can be entirely explained by local normalization distortion.
These results show that in the ongoing search for better decoding strategies for large language models, careful attention should be paid to local normalization distortion, as it has a large effect on the quality of generated text, and the size of this effect varies considerably with choice of decoding strategy. In a follow-up article, we also show how one can use local normalization distortion to detect machine-generated text.
Before diving into the technical details, it is best to first motivate the theory with a simple illustration.
2 Motivating Example
@@
Finally, having chosen the token $y_{t}$, one repeats the process to choose $y_{t+1}$ given the new context ${\bf y_{<t+1}}$. One computes the probability of a string $y_{0}\cdots y_{T}\in\mathcal{V}^{*}$ by
@@
Option 1: Global normalization
@@
Option 2: Local normalization
Renormalize conditional probabilities locally so that the probabilities of outward edges from each node sum to one. For example, we choose first word ‘a’ with probability $0.1/0.4$ and ‘the’ with probability $0.3/0.4$. With local normalization, we end up with ‘The cat sat on a fence’ being selected with probability $0.25\times 0.5=0.125$.
We see that these two different methods of normalizing probabilities yield very different probability distributions. Figure 1 uses a decoding strategy that considers the two most likely tokens at each step, known as top-k sampling with $k=2$. In the next section we formally define the array of decoding strategies we study in this article and that are used in common practice.
@@ 3 Decoding Strategies
Several newer truncation-based decoding strategies have been introduced with different (and rather clever) ways of choosing the allowed sets $\mathcal{A}_{\bf y_{<t}}$, see, for example, locally typical sampling Meister et al. (2023), $\eta$-sampling Hewitt et al. (2022) and microstat Basu et al. (2020). One observation about this body of research is that careful motivation is always given for the choice of allowed set, but little if any consideration is given to the way in which the resulting decoding strategy apportions probability mass within this allowed set.
@@ 3.2 Local Normalization Distortion
$\dfrac{1}{C}.\dfrac{1}{\prod_{i=0}^{T-t}Z_{k}({\bf y_{<t+i}})},$
@@
Given the substantial interest in implementing expensive search algorithms such as beam search to find sequences with approximately the globally maximal log probability, we present this example as initial evidence that local normalization distortion can have a substantial negative effect and is worthy of further investigation.
@@ 4 Evaluating Decoding Strategies through a Quality-Diversity Trade Off
When generating text from a language model, one may have different preferences for the ‘quality’ and ‘diversity’ of the text according to the task being performed (such as machine translation, or short story generation) Caccia et al. (2019); Wiher et al. (2022). Selecting only tokens with high likelihood produces accurate but dull text, whereas sampling from the whole distribution allows for more varied text at the risk of losing coherence. Thus, the choice of sampling algorithm affects the balance between the quality and diversity of generated text. This framing of choice of decoding strategy choice as a trade-off of quality and diversity is studied in Caccia et al. (2019); Ippolito et al. (2019); Nadeem et al. (2020); Zhang et al. (2021). In particular, in Caccia et al. (2019), the authors propose using a ‘temperature sweep’ to find a parameter $\tau$ for which temperature sampling best matches this preference. Similarly, one can adjust the parameters in top-k or nucleus sampling according to one’s preferences for diversity versus quality, since restricting token choices to the top of the distribution prioritizes quality over diversity.
@@
This notion of quality is not without its issues Meister et al. (2022), although Zhang et al. (2021) has a rather compelling graph suggesting that human judgements of quality of text are very well correlated with a continuous function of $Q(q)$.
While entropy and average log-likelihood of a distribution $q$ are imperfect, albeit frequently used, proxies for diversity and quality, they are precisely the right objects to describe mathematically the quantities maximized by the distributions resulting from top-k sampling, nucleus sampling and temperature sampling.
@@ 5 Decoding Strategies as Equilibrium States
In the last section we reviewed the literature on what a decoding strategy ought to maximize. In this section we prove results about what popular decoding strategies actually optimize. In particular, we state results of the form ‘given some context ${\bf y_{<m}}$, the probability distribution $q$ on the set $\mathcal{A}^{*}_{{\bf y_{<m}},k}$ obtained by sampling according to a certain decoding strategy is the unique probability distribution on $\mathcal{A}^{*}_{{\bf y_{<m}},k}$ which maximizes the following function…’.
In the language of ergodic theory, what we are doing is describing the outcome of a decoding strategy as an equilibrium state associated to a certain potential. While the mathematics of this section is not hard, it is useful as it allows us to ask whether the function that our decoding strategy maximizes is well aligned with the theoretical goals of a decoding strategy. In a follow-up paper we use the results of this section to theoretically justify a method of detecting language model generated text.
We use the following standard result. As usual, let $00\log 0:=$ 0.
@@
$\epsilon_{k}({{\bf w}}):=\frac{1}{\Pi_{i=0}^{T-m}Z({y_{0}\cdots y_{m}w_{m+1}\cdots w_{m+i-1}})}.$
@@
$\epsilon_{\pi}({{\bf w}}):=\frac{1}{\Pi_{i=0}^{T-m}Z({y_{0}\cdots y_{m}w_{m+1}\cdots w_{m+i-1}})}.$
@@
When nucleus or top-k sampling set $\pi<1$ or $k<|\mathcal{V}|$, they produce distributions with lower entropy than $p$, since token choice has been reduced. This process also redistributes mass from low probability tokens to higher probability tokens, reducing the average log likelihood. Thus the parameters associated with nucleus or top-k sampling allow one to prioritize quality or diversity. This is somewhat hidden in the statements of Corollaries 5.1 and 5.2, it appears only in that it is the space of distributions on $\mathcal{A}^{*}$ upon which the proxy for diversity, proxy for quality, and distortion term are maximized.
@@ 6 Experiments
All of our experiments are run using Llama 2-7B Touvron et al. (2023) on a single A100 GPU. Detailed setup for each experiment is contained in the corresponding sections and appendices.
The primary challenge to running these experiments is the computational cost of global normalization. To do this efficiently, we introduce a process based on rejection sampling in Appendix B. It is revealing that, while these computational limitations mean some experiments may only be feasibly run on short passages of generated text, this is sufficient to reliably observe that the effects predicted by our theory do indeed hold in practice.
@@ 6.1 How Large is the Distortion due to Local Normalization under Different Decoding Strategies?
Table 1: Local normalization distortion ratios over a range of comparable decoding strategies. The table is partitioned into three groups of parameters tuned so that the average amount of renormalization is equal. Reported quantiles are the absolute value of the natural log of the ratio (2).
@@
$k=5$ 1.53 3.87 8.63 14.72 20.84
$\tau=0.91$ 0.95 2.00 4.30 7.44 10.54
$p=0.77$ 0.35 1.01 2.39 4.14 6.03
$k=50$ 0.72 1.72 3.30 5.43 8.15
$\tau=0.95$ 0.47 1.11 2.44 4.01 5.74
$p=0.88$ 0.19 0.51 1.07 1.85 2.80
$k=150$ 0.34 0.97 2.02 3.50 4.97
$\tau=0.99$ 0.08 0.23 0.49 0.83 1.27
$p=0.96$ 0.06 0.13 0.27 0.47 0.70
Finding 1:
Local normalization distortion has a large effect. For example, when using temperature sampling with parameter $\tau=0.91$ to sample pairs of sequences $\bf w,\bf z$, each of length 100, we see that in half of cases, the ratio ${q_{\tau}({\bf w|c})}/{q_{\tau}({\bf z|c})}$ differs from the ratio ${p({\bf w|c})^{\frac{1}{\tau}}}/{p({\bf z|c})^{\frac{1}{\tau}}}$ by a factor of at least $\exp(4.3)\approx 74$. That is, the effect of local normalization distortion is to distort the relative probabilities of two completions by a factor of $74$.
@@ 6.2 How does Local Normalization Distortion affect the Quality-Diversity Trade Off?
[…]
Finding 3:
For both top-k and nucleus sampling, globally normalized sampling outperforms locally normalized sampling on the quality-diversity trade off. That is, at any fixed value of entropy, globally normalized sampling produces texts with higher log likelihood.
Finding 4:
As noted in Caccia et al. (2019), (locally normalized) top-k sampling performs worse than nucleus sampling on the quality-diversity trade off. However, when each sampling strategy is replaced with its globally normalized variant, the two methods perform equally well. Thus, the underperformance of top-k sampling relative to nucleus sampling is entirely explained by local normalization distortion.

### v2

@@ Abstract
Advances in hardware and language model architecture have spurred a revolution in natural language generation. However, autoregressive models compute probability distributions over next-token choices, and sampling from these distributions, known as decoding, has received significantly less attention than other design choices. Existing decoding strategies are largely based on heuristics, resulting in methods that are difficult to apply or improve in a principled manner. We develop the theory of decoding strategies for language models by expressing popular decoding algorithms as equilibrium states in the language of ergodic theory and stating the objective functions they optimize. Using this, we analyze the effect of the local normalization step required to make probabilities sum to one in top-k, nucleus, and temperature sampling. We argue that local normalization distortion is a fundamental defect of decoding strategies and quantify the size of this distortion and its effect on mathematical proxies for the quality and diversity of generated text. This yields conclusions for the design of decoding algorithms and the detection of machine-generated text.
@@ 1 Introduction
Autoregressive large-language models are poised to transform industries such as healthcare, finance and education Zhao et al. (2024). Rapid advancements in this field have been fueled by scaling Hoffmann et al. (2024), transformer network architectures Vaswani et al. (2017), and alignment processes Christiano et al. (2017); Rafailov et al. (2024). However, autoregressive language models produce conditional distributions over predicted next tokens, and these must be iteratively sampled during inference. A suite of decoding methods exist for this sampling, such as top-k, nucleus, or temperature sampling, but these are based on heuristics and our understanding of these methods is in its infancy, with most being developed through trial and error Wiher et al. (2022).
This is surprising, since the choice of decoding strategy has a profound impact on the quality of the generated text and, in some cases, may be more important than the model architecture Wiher et al. (2022). For example, greedy sampling, or the related concept of beam search, tends to produce accurate but repetitive or dull text Holtzman et al. (2020). In contrast, pure sampling produces much more varied and interesting text but can produce text which is ‘incoherent and almost unrelated to the context’ Holtzman et al. (2020).
Our primary goal is to fill this gap and initiate the theoretical study of decoding strategies. As an application of this theory, we investigate a well-known defect of popular decoding strategies, which results from truncating the distribution of possible next tokens at each step during inference. Truncation necessitates repeated renormalization to obtain valid probability distributions, and results in a phenomenon we term local normalization distortion. The fact that local normalization distorts the resulting probability distribution is well-known, but not well-understood.
Our contributions are as follows.
@@
    Develop a theoretical framework for analyzing decoding strategies. This involves describing the probability distributions $q$ produced by decoding algorithms as equilibrium states, drawing heavily on the language of ergodic theory and thermodynamic formalism Bowen (2008). We precisely state the objective function maximized by each decoding strategy; see Section 5.
@@
    Evaluate language models and decoding strategies in terms of a quality-diversity trade-off, as in Caccia et al. (2019). We show, both theoretically and empirically, that local normalization distortion negatively affects the performance of the decoding strategy. [^1] Imperfect mathematical proxies for quality and diversity of text are used in this analysis, see Section 4. Code to reproduce our experiments is available at https://github.com/TMKempton/lnd.
These results show that in the ongoing search for better decoding strategies for large language models, careful attention should be paid to local normalization distortion, as it may have a large effect on proxies for the quality of generated text, and the size of this effect varies considerably with choice of decoding strategy. Additionally, in a follow-up article Kempton et al. (2025) we show how to use local normalization distortion to detect machine-generated text.
2 A Motivating Example
@@
Having chosen the token $y_{t}$, one repeats the process to choose $y_{t+1}$ given the new context $\bf y_{<t+1}$}. One computes the probability of a string $y_{0}\cdots y_{T}\in\mathcal{V}^{*}$ by
@@
Option 1: Global normalization.
@@
Option 2: Local normalization.
Renormalize conditional probabilities locally so that the probabilities of outward edges from each node sum to one. For example, we choose the first word ‘a’ with probability $0.1/0.4$ and ‘the’ with probability $0.3/0.4$. With local normalization, we end up with ‘The cat sat on a fence’ being selected with probability $0.25\times 0.5=0.125$.
Observe that these two different methods of normalizing probabilities yield very different probability distributions.
@@ 3 Decoding Strategies
Several newer truncation-based decoding strategies have been introduced with different clever ways of choosing the allowed sets $\mathcal{A}_{\bf y_{<t}}$, see, for example, locally typical sampling Meister et al. (2023), $\eta$-sampling Hewitt et al. (2022), basis-aware truncation Finlayson et al. (2023) and microstat Basu et al. (2020). One observation about this body of research is that careful motivation is always given for the choice of allowed set, but little if any consideration is given to the way in which the resulting decoding strategy apportions probability mass within this allowed set.
@@ 3.2 Local Normalization Distortion
$\dfrac{1}{C}\cdot\dfrac{1}{\prod_{i=0}^{T-t}Z_{k}({\bf y_{<t+i}})},$
@@
There is a body of work studying global normalization in the context of constrained decoding; see, for example, Lipkin et al. (2025); Loula et al. (2025) and references therein. In these works, local normalization distortion is seen as a problem and various algorithms for sampling approximately from the globally normalized distribution are proposed.
By contrast, Gareev et al. (2024) argue that global decoding underperforms local decoding for top-k and top-$\pi$ sampling. The primary driver of the effects they observe seems to be the fact that local and global sampling produce texts of markedly different lengths (differing by a factor of four for some parameter settings), whereas we study generations of fixed length. Additionally, their approach differs from ours in that they compare quality of local and global top-k sampling at the equal values of k, whereas we compare quality of local and global top-k sampling at equal values of diversity, following the approach of Caccia et al. (2019). We justify our experimental approach in Section 6 and give further details on the difference between our work and Gareev et al. (2024) in Appendix H.
@@
Given the substantial interest in implementing expensive search algorithms such as beam search to find sequences with approximately the globally maximal log probability, we present Proposition 3.1 as initial evidence that local normalization distortion can have a substantial negative effect and is worthy of further investigation.
@@ 4 Evaluating Decoding Strategies through a Quality-Diversity Trade Off
When generating text from a language model, one may have different preferences for the ‘quality’ and ‘diversity’ of the text according to the task being performed Caccia et al. (2019); Wiher et al. (2022). By diversity, we mean the capacity for a model to produce different samples, while by quality we mean the average human judged quality of an individual sample.
This framing of choice of decoding strategy choice as a trade-off of quality and diversity is studied in Caccia et al. (2019); Ippolito et al. (2019); Nadeem et al. (2020); Zhang et al. (2021). In particular, in Caccia et al. (2019), the authors propose using a ‘temperature sweep’ to find a parameter $\tau$ for which temperature sampling best matches this preference. Similarly, one can adjust the parameters in top-k or nucleus sampling according to one’s preferences for diversity versus quality, since restricting token choices to the top of the distribution prioritizes quality over diversity.
@@
This notion of quality is not without its issues Meister et al. (2022), although Zhang et al. (2021) has a rather compelling graph suggesting that human judgements of quality of text are well correlated with $Q(q)$ except when $Q(q)$ is very high.
While entropy and average log-likelihood of a distribution $q$ are imperfect, albeit frequently used, proxies for diversity and quality, they are precisely the right objects to describe mathematically the objective functions maximized by the distributions resulting from top-k sampling, nucleus sampling and temperature sampling.
@@ 5 Decoding Strategies as Equilibrium States
In the last section we reviewed the literature on what a decoding strategy ought to maximize. In this section we prove results about what popular decoding strategies actually optimize. In particular, we state results of the form ‘given some context $\bf y_{<m}$}, the probability distribution $q$ on the set $\mathcal{A}^{*}_{{\bf y_{<m}},k}$ obtained by sampling according to a certain decoding strategy is the unique probability distribution on $\mathcal{A}^{*}_{{\bf y_{<m}},k}$ that maximizes the following objective function…’.
In the language of ergodic theory, what we are doing is describing the outcome of a decoding strategy as an equilibrium state associated to a certain potential. While the mathematics of this section is not hard, it is useful as it allows us to ask whether the function that our decoding strategy maximizes is well aligned with the theoretical goals of a decoding strategy. In a follow-up paper (Kempton et al., 2025) we apply the ideas of this section to build a novel algorithm for detecting text generated by a language model.
We use the following standard result. As usual, let $0\log 0:=0$.
@@
$\epsilon_{k}({{\bf w}}):=\frac{1}{\Pi_{i=0}^{T-m}Z({y_{0}\cdots y_{m}w_{m+1}\cdots w_{m+i-1}})},$
@@
$\epsilon_{\pi}({{\bf w}}):=\frac{1}{\Pi_{i=0}^{T-m}Z({y_{0}\cdots y_{m}w_{m+1}\cdots w_{m+i-1}})},$
@@
When nucleus or top-k sampling set $\pi<1$ or $k<|\mathcal{V}|$, they produce distributions with lower entropy than $p$, since token choice has been reduced. This process also redistributes mass from low probability tokens to higher probability tokens, reducing the average log-likelihood. Thus the parameters associated with nucleus or top-k sampling allow one to prioritize quality or diversity. This is somewhat hidden in the statements of Corollaries 5.1 and 5.2, it appears only in that it is the space of distributions on $\mathcal{A}^{*}$ upon which the proxy for diversity, proxy for quality, and distortion term are maximized.
@@ 6 Experiments
In this section we quantify the size of effects predicted in our theoretical sections. In particular, we measure the size of local normalization distortion and its effect on the quality-diversity trade off, measured through log-likelihood and entropy.
While one may, in some circumstances, care more about semantic diversity than lexical diversity, or about human judgements of quality than language model log-likelihood, effects on these metrics would be downstream of the true mathematical effect of local normalization distortion, and likely highly dependent on both the language model and the task being performed. Thus, for reasons both of stability and of quantifying our theoretical results directly, we run our experiments on metrics present in the objective functions maximized by decoding strategies.
The experiments of this section are run using Llama 2-7B Touvron et al. (2023) on a single A100 GPU. Detailed setup for each experiment is contained in the corresponding sections and appendices. In Appendix E, we repeat our experiments using Pythia 1B and 2.8B Biderman et al. (2023), and Llama 3.2 1B and 3.2 3B Grattafiori et al. (2024). As predicted above, our experimental results remain qualitatively unchanged.
The primary challenge to running these experiments is the computational cost of global normalization. To do this efficiently, we introduce a process based on rejection sampling in Appendix B.
@@ 6.1 How Large is the Distortion due to Local Normalization under Different Decoding Strategies?
Figure 2: Evaluating quality against diversity at different parameter values when decoding with top-k, nucleus sampling, and globally normalized top-k. The ranges were $k=10,11,\dots,100$ and $p=0.4,0.41,\dots,0.9$. Higher values of entropy are better, lower values of negative-$\log$-likelihood are better. We find that at any fixed value of entropy, globally normalized sampling tends to produce texts with higher log-likelihood.
Table 1: Local normalization distortion ratios over a range of comparable decoding strategies. The table is partitioned into three groups of parameters tuned so that the average amount of renormalization is approximately equal. Reported quantiles are the absolute value of the natural log of the ratio (2).
@@
$k=5$ 1.84 4.50 9.82 16.93 25.07
$\tau=0.86$ 1.18 2.76 5.97 10.80 15.65
$p=0.65$ 0.73 2.05 4.58 7.91 11.32
$k=50$ 0.60 1.43 3.26 5.60 7.92
$\tau=0.95$ 0.40 1.05 2.29 3.93 5.83
$p=0.88$ 0.18 0.44 0.95 1.76 2.70
$k=150$ 0.36 0.88 1.89 3.35 4.83
$\tau=0.98$ 0.18 0.47 0.95 1.68 2.43
$p=0.95$ 0.06 0.16 0.34 0.62 0.96
Finding 1.
Local normalization distortion has a large effect. For example, when using temperature sampling with parameter $\tau=0.86$ to sample pairs of sequences $\bf w,\bf z$, each of length 100, we see that in half of cases, the ratio ${q_{\tau}({\bf w|c})}/{q_{\tau}({\bf z|c})}$ differs from the ratio ${p({\bf w|c})^{\frac{1}{\tau}}}/{p({\bf z|c})^{\frac{1}{\tau}}}$ by a factor of at least $\exp(5.97)\approx 392$. That is, the effect of local normalization distortion is to distort the relative probabilities of two completions by a factor of $392$.
@@ 6.2 How does Local Normalization Distortion affect the Quality-Diversity Trade Off?
Given a single word context $w_{1}=\textnormal{`The '}$, we generate $30$ length $15$ samples by top-k sampling, nucleus sampling, temperature sampling and their globally normalized equivalents. We do this over a range of values of k, $\pi$ and $\tau$. For each sample $w_{2}\cdots w_{16}$, we assess the quality by computing the negative log probability $\log(p(w_{2}\cdots w_{16}|w_{1}))$. In addition, we evaluate the diversity of our sample generation process by approximating the entropy of our generation process using the Shannon-McMillan-Breimann theorem Walters (2000). That is, for each choice of decoding strategy and parameter value and each completion $w_{2}\cdots w_{16}$ generated by decoding strategy $q$, we compute $\log(q(w_{2}\cdots w_{16}|w_{1}))$. We average these values over the different completions generated for each particular decoding strategy and parameter value to approximate $H(q)$. Figure 2 shows these proxies for quality and diversity. Note that lower negative log probability and higher entropy are preferable.
Finding 3.
For both top-k and nucleus sampling, globally normalized sampling outperforms locally normalized sampling on the quality-diversity trade off. That is, at any fixed value of entropy, globally normalized sampling produces texts with higher log-likelihood.